OpenStudy (anonymous):

if f(x)=[sqrt{1-x^2}] and g(x)=[sqrt{x}] Domain and range for (f+g)of and go(f*g)=?

OpenStudy (anonymous):

hey amistre64 hello

OpenStudy (amistre64):

the domain of the other stuff still depends on the domains of the originals; see there they intersect for the overall domain

OpenStudy (amistre64):

Hi hamid

OpenStudy (amistre64):

f(x); D = {x | 1-x^2 >= 0} g(x); D = {x | x >= 0}

OpenStudy (amistre64):

the domain of the rest of the stuff would amount to the intersection of those two if both are used. and also would need to be amended if there is a division of one function by the other

OpenStudy (anonymous):

wouldn't need to be amended.it's true

OpenStudy (amistre64):

the range can then be determined by the joined pairs as they compare to the domian

OpenStudy (amistre64):

so 1-x^2 >=0 1>= x^2 +- 1 >= x ; such that x=0 is true <------------------------> [-1 0 1] and x>=0 <------------------------> [0 1 ) the intersection is defined by the interval: [0,1]

OpenStudy (amistre64):

do those [ ] mean anything important like a ceiling function perhaps?

OpenStudy (amistre64):

i think i recall [...] being notation for the greatest integer function

OpenStudy (anonymous):

i don't know

OpenStudy (anonymous):

no [..] means \[\sqrt{?}\]

OpenStudy (amistre64):

sqrt(...) tends to mean square root of some stuff so dies this imply that this is the sqrt of the sqrt of x?

OpenStudy (amistre64):

i think you might have been attempting a psuedo latex code

OpenStudy (amistre64):

\[\sqrt{1-x^2}\]

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

... ok, then my fears are allayed :)

OpenStudy (anonymous):

خیلی خری

OpenStudy (amistre64):

\[f(x)=\sqrt{1-x^2}\] \[g(x)=\sqrt{x}\] Domain and range for (f+g)of and go(f*g) \[(f(x)+g(x))\ o\ f(x)=\sqrt{1-(\sqrt{1-x^2})^2}+\sqrt{(\sqrt{1-x^2})}\] \[\] hmmm, i hope im interping that correctly

OpenStudy (anonymous):

yes yes :)D

OpenStudy (amistre64):

\[g\ o\ (f(x)*g(x))=\sqrt{\sqrt{1-x^2}*\sqrt{x}\ }\] \[g\ o\ (f(x)*g(x))=\sqrt{\sqrt{x-x^3}\ }\] \[g\ o\ (f(x)*g(x))=\sqrt[4]{x-x^3}\]

OpenStudy (amistre64):

\[(f(x)+g(x))\ o\ f(x)=\sqrt{1-1-x^2}+\sqrt[4]{1-x^2}\] \[(f(x)+g(x))\ o\ f(x)=\sqrt{-x^2}+\sqrt[4]{1-x^2}\] \[(f(x)+g(x))\ o\ f(x)=i\sqrt{x^2}+\sqrt[4]{1-x^2}\] that one might be tricky if we are to stay in the reals

OpenStudy (amistre64):

so our new function have to exist within modified domains that intersect with our elementary domain

OpenStudy (anonymous):

بله

OpenStudy (amistre64):

x-x^3 >=0 x(1-x^2)>=0 x(1-x)(1+x)>=0 defines the zeros <-----------------------> -1 0 1 test for signage within the segments 2(1-2)(1+2)>=0 + - + fails -2(1+2)(1-2)>=0 -+- pass .5(1-.5)(1+.5)>=0 +++ pass -.5(1-.5)(1+.5)>=0 -++ fail + - + - <-----------------------> -1 0 1 well, the good news is that our elementary domain is still intact :)

OpenStudy (anonymous):

well!

OpenStudy (amistre64):

for the " i sqrt(x^2)" one I think the only real value we can use is 0; so our domain would be 0 and our range 1 on that one

OpenStudy (anonymous):

فکر کنم شما خیلی زحمت کشیده اید

OpenStudy (amistre64):

\[h(x)=\sqrt[4]{x-x^3}\] we should determine of there is a max within the interval of domain for this one by taking the derivaitve \[\frac{1-3x^2}{3} \sqrt[3]{x-x^3}=0\] \[x=\pm\frac{1}{3}\] x-x^3 = 0 x(1-x^2) = 0\[x=0,\pm1\] our interests are then at x=0,1/3, and 1 h(0) = 0 h(1) = 0 \[h(1/3) = \sqrt[4]{1/3-1/27}=\sqrt{26/27}\] this looks like the highest point to me then sooo range = [0,sqrt(26/27)]

OpenStudy (amistre64):

my squiggly is a bit rusty; google translates as" you think I am too labored"

OpenStudy (anonymous):

i'm from iran where are you from? thank you very much for solve

OpenStudy (amistre64):

im sitting down here in florida

OpenStudy (amistre64):

or it might be over here in florida, my geography aint none to good

OpenStudy (anonymous):

where is florida?